OpenStudy (anonymous):

convert the width of 18in. into meters

OpenStudy (anonymous):

Approx. 0.4572

OpenStudy (amistre64):

hmm.. wed need a unit of conversion

OpenStudy (amistre64):

1 inch = .0254 meters \[\frac{18inches}{1}*\frac{.0254meters}{1inch}\] \[\frac{18\cancel{inches}}{1}*\frac{.0254meters}{1\cancel{inch}}\] \[\frac{18*.0254}{1}meters\]